Kristy has to make rental payments of $1000 at the start of every month, throughout the four-year duration of her university course. Her university fees are $4000 to be paid at the start of each year. She earns $1500 per month (paid at the end of each month) from a part-time job. Assume an interest rate of 8% p.a. and that she keeps the part-time job for the next four years. How much money, in present value terms, can she withdraw each month for the next four years?

A.

$144

B.

$126

C.

$55

D.

$177

PV value of monthly salary = PMT=-1500 N=1 FV=0 I/Y=0.67 PV=1490.02

Monthly rent                                (1000)

Monthly savings for university fee (345.65) How do I get this amount using the financial calculator?

Amount that can be withdrawn monthly 144.42
Answer #1

Solution :-

Interest rate for a year = ( 1 + 0.08 / 12 )12 - 1 = 8.304%

Present Value of College Fees paid at the starting of college = $4000 + $4000 * PVAF(8.304% , 3)

= $4000 + $4000*2.563 = $14252

Present Value of Rental Payments = $1000 + $1000 * PVAF( 0.667% , 47)

= $1000 + $1000*(34.242)

= $35242.11

Present Value of Salary = $1500 * PVAF(0.667% , 48) = $1500 * 35.0087 = $52513.05

Net Present Value of Salary + Rent + College fees = $52513.05 - $35242.11 - $14252 = $3018.94

Amount Withdrawn each month = $3018.94 / PVAF(0.667% , 48) = $3018.94 / 35.008 = $86.234

All the Answers given in it are incorrect the correct answer is $86.234